OpenStudy (anonymous):

Solve the quadratic equation: x^2+5x-36=0

Parth (parthkohli):

Use the quadratic formula here. \( \color{Black}{\Rightarrow x = \Large {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\) Do you have an idea of what a,b,c are here?

OpenStudy (anonymous):

x^2 +9x-4x-36=0 x(x+9)-4(x+9)=0 (x+9)(x-4)=0 x=-9,4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome!!!